# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ IV Discrete simulations
We avoid this instability by working in the $\tilde{m}\rightarrow 0$ limit. In Fig. S6 we numerically explore the effects of varying the dimensionless mass $\tilde{m}$ and compressive strain on the self-snapping instability of the truss, in comparison to the viscous theory developed in §III. We find that the instability is remarkably robust to both perturbations of our theory.
V Construction of our robotic metamaterials
$\tau^{a}_{i}=\begin{cases}k^{\mathrm{o}}\left(\delta\theta_{i+1}-\delta\theta_{i-1}\right),&\mathrm{if}\,\,|\delta\theta_{i+1}-\delta\theta_{i-1}|<\tau_{\mathrm{max}}/k^{\mathrm{o}}\\ \mathrm{sgn}(\delta\theta_{i+1}-\delta\theta_{i-1})\tau_{\mathrm{max}},&\mathrm{if}\,\,|\delta\theta_{i+1}-\delta\theta_{i-1}|>\tau_{\mathrm{max}}/k^{\mathrm{o}},\end{cases}$ (76)
where $\delta\theta_{i}=\theta_{i}-\theta^{0}$, $\theta^{0}$ is the rest angle and $\tau_{\mathrm{max}}$ is the torque at which the motors saturate. The active stiffness $k^{\mathrm{o}}$ was programmed by the microcontroller. The torques and angles in Eq. (76) are implemented using arbitrary digital units—bits—that are specific to our robotic implementation. We calibrate these digital values against standard physical units by measuring torque-displacement slopes for different values of the electronic feedback, as described in Ref. [32]. We find a conversion factor $846\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}{\mathrm{rad}}^{-1}\text{\,}{\mathrm{bit}}^{-1}$ taking us from digital to physical non-reciprocities $k^{\mathrm{o}}$. The coreless motor saturates at a maximum torque of $\tau_{\mathrm{max}}=$12\text{\,}\mathrm{mN}\text{\,}\mathrm{m}$\$. Each link has a length of $l_{0}=7.5\,\mathrm{cm}$. Adjacent vertices are also connected by viscoelastic hinges, whose materials properties we vary depending on the experimental protocol (§VI). For further details, see Ref. [32].
VI Experimental protocol
Figure S7: Details of our experimental buckling setup. A typical setup during buckling experiments on our robotic metamaterials, showing a chain of 12 units clamped to an air table at either end, with an Agilus 30 viscoelastic polymer scaffold attached to the top of the chain to provide hinge elasticity and dissipation. Scale bar $10\text{\,}\mathrm{cm}$.
All experiments take place on top of a custom-made low-friction air table to minimize stick-slip substrate friction. The table consists of two $1.5\times 1.5$m Plexiglas plates that sandwich 5-mm-wide air channels. The top plate is pierced by an array of holes (pitch 10 mm, diameter 1 mm) conveying air pressurized at 10 bar. Each motorized vertex floats on a thin layer of pressurized air without making contact with the table. The experiments are filmed using a Nikon D5600 camera equipped with a 50-mm lens, recording 2-Mpx images at 50 frames per second. A marker is placed on the top of each vertex and we use the Python package OpenCV to detect and track the position of each vertex with a spatial resolution of 0.8 mm. We show a typical table setup in Fig. S7.
To provide torsional rigidity and viscous dissipation to the hinges of our robotic metamaterials, we use two different viscoelastic polymers. For the data shown in Fig. 3, we use use a silicone rubber with a thickness of $2\text{\,}\mathrm{mm}$. This rubber hinge is well described with a Kelvin-Voigt model, with passive elastic torsional stiffness of $B=$20\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}{\mathrm{rad}}^{-1}$\$ and viscous dissipation $\Gamma=$0.4(1)\text{\,}\mathrm{mN}\text{\,}\mathrm{m}\text{\,}\mathrm{s}\text{\,}{\mathrm{rad}}^{-1}$\$. See Ref. [32] for calibration details. The simple viscoelastic response of this rubber means that it is well suited for the detailed calibration performed in Fig. 3. However, for longer chains (Figs. 14) we find that the dissipation of this rubber is insufficient to drive us into the viscous regime in which snapping is dominated only by the lowest modes of a longer chain.
For our chain snapping (Fig. 1) and locomotion (Fig. 4) experiments, we therefore use a highly viscous polymer. We 3D print hinges from a viscoelastic rubber-like PolyJet photopolymer (Stratasys Agilus 30). The viscoelastic response of this material has been characterised via a generalized Maxwell model in, for example, Ref. [51]. Complex viscoelastic response makes the Agilus 30 ill-suited for the detailed calibration shown in Fig. 3. However, its increased viscous dissipation renders it ideal for our locomotion experiments. In this study, we used two thicknesses of Agilus 30 bands: 3 mm and 6 mm. We determined an approximate angular stiffness $B$ and angular dissipation $\Gamma$ by placing a single robotic vertex in a rotational testing device (Instron E3000 linear-torsional) equipped with a torsional load cell (angular resolution: 2.5 arcsecs; torque resolution: 1 mN m). To determine the angular stiffness $B$ of the rubber band, the torque was measured as a function of angular deviation. The value of $B$ is taken to be the slope of the linear fit. We find $B\approx 3.90\ \mathrm{mN\cdot m/rad}$ and $B\approx\ 8.78\ \mathrm{mN\cdot m/rad}$ for the rubber band with a thickness of 3 mm and 6 mm, respectively (Fig. S8). To determine angular dissipation $\Gamma$, we place the rubber band on the linkage system, and we apply an angular perturbation to the free edge. We measure the angular displacement and extract $\Gamma$ from the decay rate. We find $\Gamma\approx 7.38\ \mathrm{mN\cdot m\cdot s/rad}$ and $\Gamma\approx 22.04\ \mathrm{mN\cdot m\cdot s/rad}$ for the rubber band with a thickness of 3 mm and 6 mm, respectively (Fig. S9). We emphasise that these measurements use a Kelvin-Voigt model for the Agilus 30, which is an oversimplification [51]. However, they indeed demonstrate the dramatically increased dissipation $\Gamma$ of the Agilus relative to a silicone rubber.
Figure S8: Calibration of angular stiffness $B$ for our Agilus-30 hinges. The restoring torque of the building block with a viscous Agilus band measured over a range of twist angles $\delta\theta$. The slope of the linear regression yields a value $B\approx 3.90\ \mathrm{mN\cdot m/rad}$ and $B\approx\ 8.78\ \mathrm{mN\cdot m/rad}$ for the rubber band with a thickness of (a) 3 mm and (b) 6 mm.
Figure S9: Calibration of angular damping coefficient $\Gamma$ for our Agilus-30 hinges. The decaying oscillations of the building block with a viscous Agilus band were fitted to the solution of a damped harmonic oscillator. We find viscous damping coefficients of $\Gamma\approx 7.38\ \mathrm{mN\cdot m\cdot s/rad}$ and $\Gamma\approx 22.04\ \mathrm{mN\cdot m\cdot s/rad}$ for the Aglius rubber band with a thickness of (a) 3 mm and (b) 6 mm.
Open chain
We assemble a 22-vertex open chain, damped using $6\text{\,}\mathrm{mm}$ Agilus-30 hinges, and loosely confined at one end. The non-reciprocity $k^{o}=34\mathrm{mN\cdot m\cdot s/rad}$, and a manual perturbation is applied to the chain’s edge, and midpoint [Fig. 2(a,b)].
Buckled chain
For the chain buckling experiments shown in Figs. 1 and  S7, we assemble an open 12-vertex filament, with each hinge damped using the $6\text{\,}\mathrm{mm}$ Agilus 30. The boundary conditions mimic a clamped-tangent boundary condition in classical buckling: At each end of the chain, the two extremal motors ($i=1,2,11,12$) are clamped in position. The boundary motors $(i=1,12)$ are clamped to external rigid bars via a custom 3D-printed pinch, which is displaced and position controlled to apply the external compression $\text{d}l$. The boundary conditions on the non-reciprocal torques mimic those implemented in our discrete theory and simulations, §III and §IV. For $i=1,12$, the active torque $\tau_{i}=0$. For $i=2$, $\tau_{2}=k^{\mathrm{o}}\delta\theta_{3}$. For $i=11$, $\tau_{11}=-k^{\mathrm{o}}\delta\theta_{10}$.
Buckled von Mises truss
[…]
Crawling experiments
Figure S10: Details of crawling experiment setup. A chain of 6 units confined via a laser cutting PMMA brace. By changing the distance between the edge unit and fastening the screws, we can maintain the chain in a buckled state. An Agilus 30 viscoelastic polymer scaffold is attached to the top of the chain to provide hinge elasticity and dissipation. Additional pieces of silicone rubber are put on the side of each unit to ensure frictional contact with the substrate. Scale bar $10\text{\,}\mathrm{cm}$.
Digging experiments
Walking experiments
Figure S11: Details of walking experiment setup. A chain of 6 units whose the top-end unit is clamped to a slide rail and the bottom-end unit can freely move. The chain is tilted to a angle $\phi$ by fastening the screw on the orange brace. An Agilus 30 viscoelastic polymer scaffold attached to the top of the chain to provide hinge elasticity and dissipation. Scale bar $10\text{\,}\mathrm{cm}$.

### v2

@@ VIII Discrete simulations
We avoid this instability by working in the $\tilde{m}\rightarrow 0$ limit. In Fig. S3 we numerically explore the effects of varying the dimensionless mass $\tilde{m}$ and compressive strain on the self-snapping instability of the truss, in comparison to the viscous theory developed in the Main Text. We find that the instability is remarkably robust to both perturbations of our theory.
IX Beam with substrate friction and follower forces
An important question is how generic is the bifurcation scenario described above. In order to address this question, we consider a beam with follower forces—also called tangential forces—which is extremely standard in the context of flagellar dynamics and active polymer literature [60, 46, 11, 22] with drag friction. In this case, the Rayleigh dissipation function is [60]
$\mathcal{R}=\frac{\eta}{2}\sum_{i=1}^{N}\dot{x}_{\text{i}}^{2}+\dot{y}_{\text{i}}^{2}+\sum_{i=2}^{N}-\vec{F^{a}_{i}}\cdot\begin{pmatrix}\dot{x}_{\text{i}}\\ \dot{y}_{\text{i}}\end{pmatrix}\text{,}$ (45)
where $\eta$ is the friction with the substrate, $\vec{F^{a}_{i}}$ is the active force on node $i$ of magnitude $F^{a}$ and aligned with the bond vector connecting node $i$ to node i1i-1. $(x_{\text{i}},y_{\text{i}})$ are the coordinates of node $i$, see Fig. S2 for a schematic. We now consider the case where nodes 1 and 4 are pinned, and express these in the same variables as defined for the odd truss in the Main Text:
$\mathcal{R}$ $=\frac{\eta}{2}\left(w_{L}\sec\theta_{L}\sin(2\theta_{L}-\theta_{R})\dot{\theta}_{L}\right)$
$\qquad+F^{a}w_{C}^{2}\left(\dot{\theta}_{L}-\dot{\theta}_{L}\right)^{2}\sec^{2}(\theta_{L}-\theta_{R})+2w_{C}w_{L}\dot{\theta}_{L}\sec(\theta_{L})\left(\dot{\theta}_{R}-\dot{\theta}_{L}\right)\cos(2\theta_{L}-\theta_{R})\sec(\theta_{L}-\theta_{R})+2w_{L}^{2}\dot{\theta}_{L}^{2}\sec^{2}(\theta_{L}),$ (46)
where we have used the same assumption as above $\dot{w}_{C}=\dot{w}_{L}=\dot{w}_{R}=0$. Using the same approach as described in the Main Text, we are able to derive the following equations of motion for the symmetric and antisymmetric angle $\theta_{S}$ and $\theta_{A}$:
$\dot{\theta}_{S}$ $=\frac{72\delta lk(3\delta l(F^{a}+48)-4(F^{a}+42))+240\delta lk^{2}-216(F^{a}+20)k}{48k}\theta_{S}$
$\qquad-\frac{3}{4}(\delta l(4(3F^{a}+k-105)-9\delta l(F^{a}-60))+9F^{a}-120)\theta_{A}+\frac{3F^{a}(113k-540)-10k^{2}+432(k+27)}{8k}\theta_{S}\theta_{A}^{2}$
$\qquad+\frac{\left(3(37F^{a}+250)k-324(F^{a}+72)+4k^{2}\right)}{4k}\theta_{A}\theta_{S}^{2}+\frac{3\left(3(41F^{a}-602)k-648(F^{a}+1)+4k^{2}\right)}{16k}\theta_{A}^{3}$
$\qquad+\frac{1}{6}(39F^{a}-10k+720)\theta_{S}^{3}$
$\dot{\theta}_{A}$ $=\left(-3(\delta l-1)(3\delta l(F^{a}+36)+F^{a}+24)-4\delta lk\right)\theta_{S}+\frac{3}{2}(-3\delta l(3\delta l-2)(F^{a}-40)+2\delta lk+3(F^{a}-20))\theta_{A}$
$\qquad+\frac{\left(-9(8F^{a}+17)k+162(F^{a}+6)+2k^{2}\right)}{2k}\theta_{A}^{2}\theta_{S}$
$\qquad-\frac{(k(21F^{a}+k+186)-3888)\theta_{A}\theta_{S}^{2}}{k}-\frac{3(27F^{a}(k-6)+(k-414)k+648)\theta_{A}^{3}}{4k}-\frac{2}{3}(6F^{a}-2k+99)\theta_{S}^{3}.$ (47)
We linearize these equations and find the Jacobian
$J=\left(\begin{array}[]{cc}\frac{1}{2}(\delta l(9\delta l(F^{a}+48)-12(F^{a}+42)+10k)-9(F^{a}+20))&\frac{3}{4}\left(9\delta l^{2}(F^{a}-60)-4\delta l(3F^{a}+k-105)-9F^{a}+120\right)\\ -3(\delta l-1)(3\delta l(F^{a}+36)+F^{a}+24)-4\delta lk&\frac{3}{2}(-3\delta l(3\delta l-2)(F^{a}-40)+2\delta lk+3(F^{a}-20))\\ \end{array}\right).$ (50)
We use this Jacobian to determine the nature of the bifurcation scenario as a function of $F^{a}$ and $\delta l$ and find a qualitatively similar Bogdanov-Takens bifurcation scenario, all organized about a critical exceptional point (Fig. S4). Hence, our results obtained in the case of non-reciprocal beam seem to pertain more generally to active filaments.
Figure S4: Existence and structure of the critical exceptional point in a buckled follower force chain. Classifying the Jacobian  (50) for the overdamped dynamics of a buckled follower force chain reveals a bifurcation scenario analogous to that of our non-reciprocal beam.
X List of Supplementary Videos
  • Video 1: Non-reciprocal buckling.
  • Video 2: Breaking reciprocity in slender filaments creates one-way flexural waves.
  • Video 3: The odd von Mises truss is a minimal model of non-reciprocal buckling.
  • Video 4: Polyfunctions of non-reciprocal filaments.
  • Video 5: Crawling and stepping are robust to perturbations.